# A discrete model for Gell-Mann matrices

### 22 Jan 2024 v1

@@ 1.3. A $3$-dimensional analogue
(52) $\lambda^{7}=\begin{pmatrix}0&0&0\cr 0&0&-i\cr 0&i&0\end{pmatrix},\quad\lambda^{8}=\frac{1}{\sqrt{3}}\begin{pmatrix}1&0&0\cr 0&1&0\cr 0&0&-2\end{pmatrix}$
@@ 2.1. The binary tetrahedral group
At the more abstract group theory level, we have constructed a semi-direct product of $Q_{8}$ by $Z_{3}$, which we denote $Q_{8}\rtimes Z_{3}$. The subgroup $Q_{8}$ generated by the matrices $K$, $J$, $I$ in (25) is normal, but the subgroup $Z_{3}$ generated by the above matrix $W$ is not. In place of the relations $IW=WI$, $JW=WJ$ that hold in the direct product, we have the relations $IW=WJ$ and $JW=WIJ$. The continuous analogue of $Q_{8}$ is $SU(2)$, and the continuous analogue of $Z_{3}$ is $U(1)$, so that the effect of this semi-direct product is to create a more complicated relationship between $SU(2)$ and $U(1)$ than the direct product $SU(2)\times U(1)$. In physics language this relationship is a ‘mixing’ of $SU(2)$ with $U(1)$ at the quantum level.
@@ 2.2. A ternary representation
(98) $\frac{t}{3}\begin{pmatrix}1&1&1\cr 1&v&w\cr 1&w&v\end{pmatrix},\quad\frac{t}{3}\begin{pmatrix}1&v&v\cr w&v&w\cr w&w&v\end{pmatrix},$
@@
At this stage we have a group of order $216$ that combines finite analogues of $SU(3)$ and $SU(2)$ into a single group. This is a semi-direct product $G_{27}\rtimes Q_{8}$, in which the subgroup $G_{27}$ is normal, but the subgroup $Q_{8}$ is not. Since it is not a direct product, it does not give rise to a direct product $SU(3)\times SU(2)$, but instead ‘mixes’ the two gauge groups.
@@ 2.3. A complex reflection group
If we expand the complex $1$-spaces to real $2$-spaces, then the six complex scalar multiples of $12$ complex $3$-vectors become two real scalar multiples of $36$ real $6$-vectors, and the corresponding $36$ real reflections generate the Weyl group of $E_{6}$. In other words, the group $G_{27}\rtimes Q_{8}\rtimes Z_{3}$ is a subgroup of the Weyl group of $E_{6}$, and is in fact a maximal subgroup of the rotation part of the Weyl group. It is possible that this fact may have led to the appearance of $E_{6}$ symmetry in certain Grand Unified Theories (GUTs) such as [14, 15].
@@ 3.1. Continuous gauge groups
So far we have only looked at the combinatorial structure of the group, which gives a qualitative but not quantitative picture of the underlying physics. In order to introduce measurable physical quantities such as mass, momentum and energy, and allocate them to elementary (or composite) particles, we need to introduce real (or complex or quaternionic) representations, and use real (or complex or quaternionic) gauge groups to choose coordinates for the representations. There are three fundamental representations, described above, given by the embeddings of finite groups in Lie groups as follows:
@@
In the case of $U(1)$ acting on a complex $1$-dimensional representation, there is one physical parameter to measure, represented as an angle, or phase, in the complex plane. In the case of $SU(2)$ acting on a $1$-dimensional quaternionic representation, there are three such angles. Therefore by studying the representations of the binary tetrahedral group $Q_{8}\rtimes Z_{3}$, we might expect to find four of the fundamental mixing angles of the Standard Model, including the electro-weak mixing angle (Weinberg angle). A tentative identification of these four angles is obtained in [9].
@@ 3.2. Clifford theory
At the fourth stage, we have the fundamental pseudoreal representation as quaternions, with $Z_{3}$ as the corresponding ‘inertial quotient’, so that the other representations at this stage are obtained by tensoring with the representations of $Z_{3}$. This gives us a pair of complex conjugate two-dimensional representations, or equivalently a real four-dimensional representation. Putting all these stages together gives us the representations of the binary tetrahedral group, which are very well known and have been widely studied. The character table can be found in [21, p.404] and is reproduced in Table 1 for convenience, in a shorthand form in which conjugacy classes that are scalar multiples of each other are grouped together:
@@ 3.3. The Hessian group
Coming now to $G_{27}$, whose character table can be found in [21, p.400] we first take a $1$-dimensional complex representation, with inertial quotient $Z_{3}$, and induce up to the whole group $G_{27}\rtimes Q_{8}\rtimes Z_{3}$, to get an $8$-dimensional real representation, together with its tensor products with the representations of $Z_{3}$. The scalars in $G_{27}$ act trivially on these representations, so they are representations of a group $(Z_{3}\times Z_{3})\rtimes Q_{8}\rtimes Z_{3}$ of order $216$, which has been known as the Hessian group since the 1870s.
@@ 3.4. The faithful characters
Table 4. Abbreviated table of characters for quarks
It is not really necessary to write out this table in full, since, as noted above, the seven rows are obtained by multiplying the first row with the seven irreducible characters of the binary tetrahedral group, that is the first seven characters of Table 2. If we take the latter to describe properties of leptons, then Table 4 describes the corresponding properties of quarks. Indeed, if we have the representations $\mathbf{1}b$, $\mathbf{2}a$ and $\mathbf{3}b$ then we can calculate all the other representations easily as tensor products, subject only to the relations
@@
Hence it is only necessary to gauge the representations $\mathbf{1}a$, $\mathbf{2}a$ and $\mathbf{3}b$, using the gauge groups $U(1)$, $SU(2)$ and $SU(3)$ respectively.
@@ 4.1. The electroweak gauge
The $U(1)$ gauge group contains only one parameter, or ‘mixing angle’, that identifies the directions of the real and imaginary parts of the representation. This angle has been computed in [9] as $33.024^{\circ}$. The method of calculation was to superimpose the mass axis on an equilateral triangle marked with the three generations of electrons on the vertices. Hence it is expected that this mixing angle is one of the lepton mixing angles in the Standard Model.
The $SU(2)$ gauge group similarly contains exactly three independent parameters. The angles calculated in [9] were $28.165516^{\circ}$, $66.7276^{\circ}$ and $2.337325^{\circ}$. The method here was to identify discrete quaternions appropriate for four particles (the three generations of electron, plus a proton) that are ‘fundamental’ as far as electroweak interactions are concerned, and again superimpose a mass axis. We should certainly expect one of these angles to be the electroweak mixing angle, but it is not clear a priori whether the other two are lepton-mixing (PMNS) or quark-mixing (CKM) angles. However, if we take into account that the embedding of $Q_{8}\rtimes Z_{3}$ into $SU(3)$ is described by the representation $\mathbf{1}b+\mathbf{2}c$, we should expect one ‘scalar’ (i.e. CP-violating phase) and one generation-mixing angle.
@@ 4.3. Identifying particles
However it is done, we must start by identifying the particles we use with certain vectors or subspaces in $\mathbb{C}^{3}$. Therefore we must begin by examining the geometry of this $3$-dimensional space, using the complex reflection group as a guide. The most important vectors are those that lie in the $12$ ‘mirrors’. The particular scalar multiple chosen is irrelevant for defining the reflection, so to begin with at least we can ignore the scalar multiplication. The Hessian group then acts on this set of $12$ one-spaces in four triplets, as shown in (123). The reflections in any mirror fix the mirrors in its column, and cycle the other three columns. The normal subgroup $G_{27}/Z_{3}$ of order $9$ contains four subgroups of order $3$, each of which fixes the mirrors in one column, and cycles the mirrors in the other three columns (not always in the same direction!). It is therefore tempting to identify these $12$ mirrors with the $12$ types of elementary fermions, and use the scalars for unobservable properties such as colours of quarks and spin direction of electrons. However, the details of such an identification are tricky.
@@ 5.1. Finite fields
The $3$-dimensional space over $\mathbb{F}_{4}$ on which $U(3,\mathbb{F}_{4})$ acts contains $634^{3}-1=$ 63 vectors, each in a set of $3$ scalar multiples, and therefore forming $63/3=21$ one-dimensional subspaces. These $21$ ‘points’ of the geometry divide into $12$ ’odd’ (non-singular) and $9$ ’even’ (singular) points. Each point is perpendicular to a ‘line’. The $12$ non-singular lines each contain $2$ non-singular and $3$ singular points, while the $9$ singular lines each contain $1$ singular point and $4$ non-singular points.
@@ 5.2. Interpretations and speculations
In other words, the proposed splitting of the ‘generation’ symmetry into two separate triplet symmetries is strongly supported by experiment. But there is a crucial difference between the electron generation symmetry, in which the mass changes very obviously by several orders of magnitude between generations, and the neutrino generation symmetry, in which no mass can be detected directly at all, but is instead inferred from the details of the theoretical model.
It is reasonable to suppose, therefore, that the electron generation symmetry acts to change the mass of the quarks, while the neutrino generation symmetry has no detectable effect on quark mass. In other words, the neutrino generation symmetry effectively acts on the colours of the quarks rather than their generation. But this implies that the neutrino generation should rather be interpreted as a colour, and that neutrino oscillation should be interpreted as a change in colour, and must therefore be mediated by gluons.

### 25 Jan 2024 v2

@@ 1.1. Context, aims and objectives
A secondary aim is to show how the (modified) Gell-Mann matrices can act on an extended (quaternionic) version of the Dirac matrices, and thereby change the local shape of spacetime. We then investigate (at a qualitative level) the possibility of using this action to develop a generally covariant model of quantum gravity, that appears to reduce to General Relativity [13, 14, 15] if the three generations of electrons are ignored, and all matter is assumed to consist of neutrons.
@@ 1.3. A $3$-dimensional analogue
(52) $\lambda^{7}=\begin{pmatrix}0&0&0\cr 0&0&-i\cr 0&i&0\end{pmatrix},\quad\lambda^{8}=\frac{1}{\sqrt{3}}\begin{pmatrix}1&0&0\cr 0&1&0\cr 0&0&-2\end{pmatrix}.$
@@ 2.1. The binary tetrahedral group
At the more abstract group theory level, we have constructed a semi-direct product of $Q_{8}$ by $Z_{3}$, which we denote $Q_{8}\rtimes Z_{3}$. The subgroup $Q_{8}$ generated by the matrices $K$, $J$, $I$ in (25) is normal, but the subgroup $Z_{3}$ generated by the above matrix $W$ is not. In place of the relations
(92) $IZ=ZI,\quad JZ=ZJ$
that define the direct product, $Q_{8}\times Z_{3}=\langle I,J\rangle\times\langle Z\rangle$, we have the relations
(93) $IW=WJ,\quad JW=WIJ$
in the semidirect product $Q_{8}\rtimes Z_{3}=\langle I,J\rangle\rtimes\langle W\rangle$.
The continuous analogue of $Q_{8}$ is $SU(2)$, and the continuous analogue of $Z_{3}$ is $U(1)$, so that the effect of this semi-direct product is to create a more complicated relationship between $SU(2)$ and $U(1)$ than the direct product $SU(2)\times U(1)$. In physics language this relationship is a ‘mixing’ of $SU(2)$ with $U(1)$ at the quantum level. In the case of electro-weak mixing, $Z$ represents a generator for the group $U(1)_{Y}$ of weak hypercharge, while $W$ represents a generator of the group $U(1)_{em}$ of (electromagnetic) charge, and $I,J,K$ represent the three components of weak isospin, as generators for $SU(2)_{L}$.
@@ 2.2. A ternary representation
(100) $-\frac{t}{3}\begin{pmatrix}1&1&1\cr 1&v&w\cr 1&w&v\end{pmatrix},\quad-\frac{t}{3}\begin{pmatrix}1&v&v\cr w&v&w\cr w&w&v\end{pmatrix},$
@@
At this stage we have a group of order $216$ that combines finite analogues of $SU(3)$ and $SU(2)$ into a single group. This is a semi-direct product $G_{27}\rtimes Q_{8}$, in which the subgroup $G_{27}$ is normal, but the subgroup $Q_{8}$ is not. Since it is not a direct product, it does not give rise to a direct product $SU(3)\times SU(2)$, but instead ‘mixes’ the two gauge groups. If we compute the action of the matrices (100) on the 3-vectors, we find that $(0,1,-1)$ is an eigenvector, with eigenvalue $1$. Hence $Q_{8}$ embeds in a subgroup $SU(2)$ of $SU(3)$, which gives the required mixing.
@@ 2.3. A complex reflection group
If we expand the complex $1$-spaces to real $2$-spaces, then the six complex scalar multiples of $12$ complex $3$-vectors become two real scalar multiples of $36$ real $6$-vectors, and the corresponding $36$ real reflections generate the Weyl group of $E_{6}$. In other words, the group $G_{27}\rtimes Q_{8}\rtimes Z_{3}$ is a subgroup of the Weyl group of $E_{6}$, and is in fact a maximal subgroup of the rotation part of the Weyl group. It is possible that this fact may have led to the appearance of $E_{6}$ symmetry in certain Grand Unified Theories (GUTs) such as [17, 18]. Unfortunately, however, the loss of the unitary structure of the $3$-dimensional representation appears to result in a loss of contact to real world physics. For this reason we are not considering models of $E_{6}$ type in this paper, but sticking rigorously to models that can be justified by experiment.
@@ 3.1. Continuous gauge groups
So far we have only looked at the combinatorial structure of the group, which gives a qualitative but not quantitative picture of the underlying physics. In order to introduce measurable physical quantities such as mass, momentum and energy, and allocate them to elementary (or composite) particles we need to introduce real (or complex or quaternionic) representations, and use real (or complex or quaternionic) gauge groups to choose coordinates for the representations.
There are three fundamental representations, described above, given by the embeddings of finite groups in Lie groups as follows:
@@
In the case of $U(1)$ acting on a complex $1$-dimensional representation, we are left therefore with one physical parameter to measure, which can be represented as an angle, or phase, in the complex plane. In the case of $SU(2)$ acting on a $1$-dimensional quaternionic representation, there are three such angles. Therefore by studying the representations of the binary tetrahedral group $Q_{8}\rtimes Z_{3}$, we might expect to find four of the fundamental mixing angles of the Standard Model, including the electro-weak mixing angle (Weinberg angle). A tentative identification of these four angles is obtained in [9].
@@ 3.2. Clifford theory
At the fourth stage, we have the fundamental pseudoreal representation as quaternions, with $Z_{3}$ as the corresponding ‘inertial quotient’, so that the other representations at this stage are obtained by tensoring with the representations of $Z_{3}$. This gives us a pair of complex conjugate two-dimensional representations, or equivalently a real four-dimensional representation. Putting all these stages together gives us the representations of the binary tetrahedral group, which are very well known and have been widely studied. The character table can be found in [24, p.404] and is reproduced in Table 1 for convenience, in a shorthand form in which conjugacy classes that are scalar multiples of each other are grouped together:
@@ 3.3. The Hessian group
Coming now to $G_{27}$, whose character table can be found in [24, p.400] we first take a $1$-dimensional complex representation, with inertial quotient $Z_{3}$, and induce up to the whole group $G_{27}\rtimes Q_{8}\rtimes Z_{3}$, to get an $8$-dimensional real representation, together with its tensor products with the representations of $Z_{3}$. The scalars in $G_{27}$ act trivially on these representations, so they are representations of a group $(Z_{3}\times Z_{3})\rtimes Q_{8}\rtimes Z_{3}$ of order $216$, which has been known as the Hessian group since the 1870s.
@@
From this character table one can easily calculate the anti-symmetric square of $\mathbf{8}a$, and we find
(130) $\Lambda^{2}(\mathbf{8}a)=\mathbf{2}b+\mathbf{2}c+\mathbf{8}a+\mathbf{8}b+\mathbf{8}c.$
In particular, there is a unique anti-symmetric (anti-commutative) algebra structure on $\mathbf{8}a$, namely the Lie algebra. The symmetric square is similarly
(131) $S^{2}(\mathbf{8}a)=\mathbf{1}a+\mathbf{3}a+\mathbf{8}a+\mathbf{8}a+\mathbf{8}b+\mathbf{8}c.$
Hence there are two independent symmetric (commutative) products on $\mathbf{8}a$ as well. One can therefore construct a wide variety of different algebra structures on $\mathbf{8}a$, some of which have been studied in the physical literature [25, 26].
Indeed, the corresponding representations of the compact Lie group $SU(3)$ are as follows:
(132) $\Lambda^{2}(\mathbf{8})$ $=\mathbf{8}+\mathbf{10}+\overline{\mathbf{10}}$
(133) $S^{2}(\mathbf{8})$ $=\mathbf{1}+\mathbf{8}+\mathbf{27}$
so that there is a (unique up to scalars) commutative algebra invariant under $SU(3)$, plus an infinite family of algebras that are neither commutative nor anti-commutative.
Table 4. Abbreviated table of characters for quarks
@@ 3.4. The faithful characters
It is not really necessary to write out this table in full, since, as noted above, the seven rows are obtained by multiplying the first row with the seven irreducible characters of the binary tetrahedral group, that is the first seven characters of Table 2. If we take the latter to describe properties of leptons, then Table 4 describes the corresponding properties of quarks.
Indeed, if we have the representations $\mathbf{1}b$, $\mathbf{2}a$ and $\mathbf{3}b$ then we can calculate all the other representations easily as tensor products, subject only to the relations
@@
Hence it is only necessary to gauge the representations $\mathbf{1}b$, $\mathbf{2}a$ and $\mathbf{3}b$, using the gauge groups $U(1)$, $SU(2)$ and $SU(3)$ respectively.
@@ 4.1. The electroweak gauge
The $U(1)$ gauge group contains only one parameter, or ‘mixing angle’, that identifies the directions of the real and imaginary parts of the representation. This angle has been computed in [9] as $33.024^{\circ}$. The method of calculation was to superimpose the mass axis on an equilateral triangle marked with the three generations of electrons on the vertices. Hence it is expected that this mixing angle is one of the lepton mixing angles in the Standard Model. However, it should be noted that this is an angle between the mass axis and one of the edges of the equilateral triangle, and it is not clear a priori which edge of the triangle to use, or in which direction. Of the six possible angles, only this one is consistent with the experimental values of the mixing angles. Hence we are not making a prediction here, but fitting the model to the experimental data, using a discrete fitting parameter of $60^{\circ}$.
The $SU(2)$ gauge group similarly contains exactly three independent parameters. The angles calculated in [9] were $28.165516^{\circ}$, $66.7276^{\circ}$ and $2.337325^{\circ}$. The method here was to identify discrete quaternions appropriate for four particles (the three generations of electron, plus a proton) that are ‘fundamental’ as far as electroweak interactions are concerned, and again superimpose a mass axis. There is again a discrete set of possible angles to measure, and again we are fitting the model to the experimental data rather than making a specific prediction.
We should certainly expect one of these angles to be the electroweak mixing angle, but it is not clear a priori whether the other two are lepton-mixing (PMNS) or quark-mixing (CKM) angles. However, if we take into account that the embedding of $Q_{8}\rtimes Z_{3}$ into $SU(3)$ is described by the representation $\mathbf{1}b+\mathbf{2}c$, we should expect one ‘scalar’ (i.e. CP-violating phase) and one generation-mixing angle.
@@ 4.3. Identifying particles
However it is done, we must start by identifying the particles we use with certain vectors or subspaces in $\mathbb{C}^{3}$. Therefore we must begin by examining the geometry of this $3$-dimensional space, using the complex reflection group as a guide. The most important vectors are those that lie in the $12$ ‘mirrors’. The particular scalar multiple chosen is irrelevant for defining the reflection, so to begin with at least we can ignore the scalar multiplication. The Hessian group then acts on this set of $12$ one-spaces in four triplets, as shown in (125).
The reflections in any mirror fix the mirrors in its column, and cycle the other three columns. The normal subgroup $G_{27}/Z_{3}$ of order $9$ contains four subgroups of order $3$, each of which fixes the mirrors in one column, and cycles the mirrors in the other three columns (not always in the same direction!). It is therefore tempting to identify these $12$ mirrors with the $12$ types of elementary fermions, and use the scalars for unobservable properties such as colours of quarks and spin direction of electrons. However, the details of such an identification are tricky.
@@ 5.1. Finite fields
The $3$-dimensional space over $\mathbb{F}_{4}$ on which $U(3,\mathbb{F}_{4})$ acts contains $634^{3}-1=$ 63 vectors, each in a set of $3$ scalar multiples, and therefore forming $63/3=21$ one-dimensional subspaces. These $21$ ‘points’ of the geometry divide into $12$ ‘odd’ (non-singular) and $9$ ‘even’ (singular) points. Each point is perpendicular to a ‘line’. The $12$ non-singular lines each contain $2$ non-singular and $3$ singular points, while the $9$ singular lines each contain $1$ singular point and $4$ non-singular points.
@@ 5.2. Interpretations and speculations
Thus the proposed splitting of the ‘generation’ symmetry into two separate triplet symmetries is supported by experiment, at least to a certain extent. But there is a crucial difference between the electron generation symmetry, in which the mass changes very obviously by several orders of magnitude between generations, and the neutrino generation symmetry, in which no mass can be detected directly at all, but is instead inferred from the details of the theoretical model. It is reasonable to suppose, therefore, that the electron generation symmetry acts to change the mass of the quarks, while the neutrino generation symmetry has no detectable effect on quark mass.